What is the probability of drawing 3 cards without replacement from a deck of cards and getting three kings




1
Expert's answer
2021-11-12T15:02:06-0500

Let us use the standard desk of 52 card.

"P = \frac{4}{52} \times \frac{3}{51} \times \frac{2}{50} = \frac{24}{132600} = 1.809 \times 10^{-4}"

"\frac{4}{52}" the probability of drawing one from 4 kings

"\frac{3}{51}" the probability of drawing one from 3 kings and remaining 51 cards

"\frac{2}{50}" the probability of drawing one from 2 kings and remaining 50 cards